OpenStudy (anonymous):

What is 0.28 written as fraction in simplest form?

OpenStudy (anonymous):

\[\frac{ 28 }{ 100 }\]

OpenStudy (anonymous):

\[0.28=28\times 10^{-2}\] \[28\times 10^{-2}=\frac{28}{10^{2}}\] \[\frac{28}{10^{2}}=\frac{28}{100}\]

OpenStudy (7thgrader):

@darthtony and @lovemenplayhard you guys are stupid as fu.ck it says in simpilist form

OpenStudy (7thgrader):

@mei66 i know it's lat but the correct answer was 7/25

OpenStudy (7thgrader):

ok two years late but still